Solve the problems below and show the complete solution.

A vertical spring has a force constant k of 130N/m and a mass of 0.645kg is placed on top of it.
a) Calculate the mass of the object in N. F = w g
b) Calculate the x if the spring be compressed. F = k x
x = F / k
1
Expert's answer
2020-12-08T10:52:36-0500

a)


"F=mg=(0.645)(9.8)=6.32\ N"

b)


"x=\frac{F}{k}=\frac{6.32}{130}=0.0486\ m"